Fermi National Laboratory
 

Accelerator Update for October 2, 2006 - October 6, 2006

 

Wednesday October 3

The day shift began with Operations monitoring the Recycler stash (26.0E10) and the Antiproton Source (Pbar) stack (30.4E10), with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the colliding beam, and with beam in Linac, but with no beam to Booster or the Main Injector (MI) due to a Booster access.

Operations began turning on the Booster at 8:33 AM.

The Recycler completed stashing antiprotons at 8:49 AM. Pbar resumed stacking and MiniBooNE resumed taking beam.

At 9:42 AM, Operations tried to inject protons into the Tevatron (TeV), but a power supply didn’t ramp (T:ILAM) and a kicker (F17K) didn’t fire. TeV experts investigated.

Operations successfully circulated beam in the TeV at 10:10 AM.

Operations prepared D0 for an access at 10:21 AM. They came out of access at 11:55 AM.

Booster experts began a short study period at 1:11 PM.

Operations prepared the TeV for a dry squeeze at 2:17 PM.

The Recycler completed a stash at 2:24 PM.

Operations prepared D0 for access at 6:10 PM.

The Recycler completed a stash at 6:52 PM.

Operations put the TeV into shot setup at 7:46 PM.

At 9:21 PM, the NuMI horn tripped off. It reset okay.

At 10:03 PM, a MI power supply (I:Q703) tripped off and dropped the beam permits for MI, the P1 transfer line, and NuMI.

Operations established store 5002 at 10:32 PM with an initial average luminosity of 127.04E30.

The Recycler completed a stash at 11:40 PM.

Thursday October 4

The midnight shift began with Operations monitoring store 5002 (Luminosity 92.67E30) and the stash (71.1E10), with Pbar stacking (21.3E10), and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash at 3 AM.

The Recycler completed a stash at 6:27 AM.

Operations reported at 8:45 AM that they set a new stacking record over the midnight shift with a rate of 20.5 mA per hour.

A MI expert began a four-hour slipstacking study at 10:28 AM.

The Recycler completed a stash around 2:24 PM.

The NuMI horn tripped off at 5:56 PM on a ground fault indication. It reset okay.

Tevatron Electron Lens (TEL) experts began a study at 8:19 PM.

The Recycler completed a stash at 8:19 PM.

Operations terminated store 5002 at 9 PM.

Operations put the TeV into shot setup at 9:37 PM.

Operations established store 5003 at 11:52 PM with an initial average luminosity of 209.13E30.

Friday October 5

The midnight shift began with Operations monitoring store 5003 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Around 12:20 AM, store 5003 aborted due to a TeV sector B3 VCB opening. It opened due to an oil level fault.

At 1:38 AM, a power supply expert reported that the B3 oil level was okay. He reported that the gauge was bad and jumpered it out.

Operations had the TeV ramping by 2:37 AM, but five minutes later the ramp dumped due to a TeV sector D2 lead failure.

Operations had the TeV ramping again by 3:07 AM. Operations intentionally dumped the ramp at 3:54 AM to change out a D2 VFC card.

Operations put the TeV through dry squeeze at 5:28 AM.

The Recycler completed a stash at 6:16 AM.

Operations put the TeV through another dry squeeze at 7:42 AM.

The Plans for Friday and the Weekend

The plans for today are to stack, stash, and store. The Run Coordinator may allow Pbar experts to conduct some studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/06/2006    email Fermilab